# Schur-Convex Curvature on Dihedral Exponential Families and the Golden-Ratio Stationary Point

### v1

@@ Appendix A Modular identities for the golden-ratio lock-in
This appendix provides the cancellation underlying Proposition 5.1. The key observation is that at $q_{\star}=\varphi^{-2}=(3-\sqrt{5})/2$, the parameter satisfies the quadratic identity
@@
the minimal polynomial of $\varphi^{-2}$ over $\mathbb{Q}$.
@@ A.3 Three-cycle identity at $q_{\star}$
When $N\equiv 0\pmod{3}$, the dihedral action partitions the indices $s\in\{1,\dots,N\}$ into three orbits modulo $3$. Evaluating at $q_{\star}$ and using (8) to reduce every $q^{m}$ to an affine form in $\{1,q_{\star}\}$, each of the sums $S_{k}(q_{\star})$ and their $\theta$-derivatives decompose into contributions from the three orbits with the same two-dimensional basis $\{1,q_{\star}\}$.
@@
Express $I_{1}^{\prime}(\theta)=I_{2}-I_{1}^{2}$ and $I_{2}^{\prime}(\theta)=I_{3}-I_{1}I_{2}$ using (9). Write $S_{k}(q_{\star})=\alpha_{k}+\beta_{k}q_{\star}$ by reducing all powers via (8). Because the $N$ terms split evenly across the three congruence classes modulo $3$ when $N\equiv 0\pmod{3}$, the numerators and denominators of $I_{1}^{\prime},I_{2}^{\prime}$ reduce to the same $\{1,q_{\star}\}$-basis. Comparing coefficients gives a linear relation of the form (10), with $\Lambda(N)$ a rational function of $N$ and the $\alpha_{k},\beta_{k}$. (An explicit evaluation for $N=12$ is carried out in Appendix C.) ∎
@@ A.4 Cancellation in $F^{\prime}_{\mathrm{red}}$
$F^{\prime}_{\mathrm{red}}(\theta)=-\tfrac{8}{Nm_{\rho}^{2}}I_{1}I_{1}^{\prime}+\tfrac{1}{N}\big(B\,I_{2}^{\prime}+(2A-2B)\,I_{1}I_{1}^{\prime}\big).$
Evaluating at $\theta_{\star}=\ln q_{\star}$ and using Lemma A.1 yields
$F^{\prime}_{\mathrm{red}}(\theta_{\star})=\frac{1}{N}\Big(\,B\,\Lambda(N)+2A-2B-\frac{8}{m_{\rho}^{2}}\,\Big)\,I_{1}(\theta_{\star})\,I_{1}^{\prime}(\theta_{\star}).$ (11)
For the $D_{12}$ case with $m_{\rho}^{2}=2$, the projector metric fixes $A,B$ (Appendix G) and one checks the identity
@@
so the bracket in (11) vanishes and hence $F^{\prime}_{\mathrm{red}}(\theta_{\star})=0$. This is the $N=12$ instance of Proposition 5.1.
@@ A.5 Role of convexity
By Theorem 3.1, $\kappa_{\mathrm{Schur}}(\theta)$ is convex; under the strict-convexity criterion (Lemma 4.4) the stationary point is unique (Theorem 5.2). Thus $q_{\star}=\varphi^{-2}$ is not only a stationary point but the unique lock-in of the system.
Summary. The minimal polynomial (8) forces all folded sums to collapse onto the two-dimensional basis $\{1,q_{\star}\}$. When $N\equiv 0\pmod{3}$, the dihedral three-orbit structure yields the proportionality (10), and the bracket in (11) cancels exactly at $N=12$ by (12). This proves Proposition 5.1.
@@ Appendix D Embedding and strong convergence criterion
Let $\mathcal{E}_{N}$ be the span of exponentials $e^{ik\theta}$ with $0<|k|\leq N/2-1$. The unitary $U_{N}:\mathbb{C}^{N}\to\mathcal{E}_{N}$ maps discrete Fourier coefficients to trigonometric polynomials. For $L_{N}$ circulant, the lifted operator $T_{N}=U_{N}L_{N}U_{N}^{\ast}$ acts as a Fourier multiplier with discrete symbol $m_{N}$ [3].
@@
If $T_{N}$ are uniformly bounded ($\sup_{N}\|T_{N}\|<\infty$) and $T_{N}f\to Tf$ for all $f$ in a dense subset of $L^{2}_{0}(\mathbb{S}^{1})$ (e.g. trigonometric polynomials), then $T_{N}\to T$ strongly on $L^{2}_{0}(\mathbb{S}^{1})$ [6].

### v2

@@ Appendix A Modular identities for the golden-ratio lock-in
This appendix records the modular reductions used in the proof of Proposition 5.1. At the stationary candidate $q_{\star}=\varphi^{-2}=(3-\sqrt{5})/2$, the parameter satisfies the quadratic identity
@@
the minimal polynomial of $\varphi^{-2}$ over $\mathbb{Q}$. When $N\equiv 0\pmod{12}$, the $D_{N}$ action decomposes the nonzero residues into $12$-cycles, and the folded moments $I_{m}(\theta)$ inherit reduction rules consistent with (9). These identities are what allow $I_{2}^{\prime}(\theta_{\star})$ and $I_{1}(\theta_{\star})I_{1}^{\prime}(\theta_{\star})$ to satisfy the cancellation appearing in the proof.
@@ A.3 Three-cycle identity at $q_{\star}$
When $N\equiv 0\pmod{3}$, the dihedral action partitions the indices $s\in\{1,\dots,N\}$ into three orbits modulo $3$. Evaluating at $q_{\star}$ and reducing each power of $q_{\star}$ via (9), every folded sum $S_{k}(q_{\star})$ and its $\theta$-derivative collapses to an affine combination of $\{1,q_{\star}\}$. Thus the three-orbit decomposition forces the folded moments to lie in a common two-dimensional space.
@@
Using (10), $I_{1}^{\prime}(\theta)=I_{2}-I_{1}^{2}$ and $I_{2}^{\prime}(\theta)=I_{3}-I_{1}I_{2}$. Write $S_{k}(q_{\star})=\alpha_{k}+\beta_{k}q_{\star}$ after reducing powers using (9). For $N\equiv 0\pmod{3}$, each of the three congruence classes contributes equally, so the numerators and denominators of $I_{1}^{\prime}$ and $I_{2}^{\prime}$ have the same $\{1,q_{\star}\}$-basis. Hence the ratio $I_{2}^{\prime}(\theta_{\star})/I_{1}^{\prime}(\theta_{\star})$ is a scalar, giving (11). An explicit calculation for $N=12$ appears in Appendix C. ∎
@@ A.4 Cancellation in $F^{\prime}_{\mathrm{red}}$
$F^{\prime}_{\mathrm{red}}(\theta)=-\tfrac{8}{Nm_{\rho}^{2}}I_{1}I_{1}^{\prime}+\tfrac{1}{N}\bigl(B\,I_{2}^{\prime}+(2A-2B)\,I_{1}I_{1}^{\prime}\bigr).$
Evaluating at $\theta_{\star}=\ln q_{\star}$ and applying Lemma A.1 gives
$F^{\prime}_{\mathrm{red}}(\theta_{\star})=\frac{1}{N}\Bigl(B\,\Lambda(N)+2A-2B-\tfrac{8}{m_{\rho}^{2}}\Bigr)I_{1}(\theta_{\star})\,I_{1}^{\prime}(\theta_{\star}).$ (12)
For $D_{12}$ with $m_{\rho}^{2}=2$, the projector geometry determines $A,B$ (Appendix G). Direct evaluation shows
@@
so the bracket in (12) vanishes identically. Thus $F^{\prime}_{\mathrm{red}}(\theta_{\star})=0$, the $N=12$ instance of Proposition 5.1.
@@ A.5 Role of convexity
By Theorem 3.1, $\kappa_{\mathrm{Schur}}(\theta)$ is convex. Under the strict-convexity criterion (Lemma 4.4), the stationary point is unique. Hence $q_{\star}=\varphi^{-2}$ is the unique lock-in of the system.
Summary. The minimal polynomial (9) collapses all folded sums onto the basis $\{1,q_{\star}\}$. For $N\equiv 0\pmod{3}$, the three-cycle decomposition gives the proportionality (11), and in the dihedral case $N=12$, the identity (13) forces the bracket in (12) to vanish. This completes the proof of Proposition 5.1.
@@ Appendix D Embedding and strong convergence criterion
Let $\mathcal{E}_{N}$ be the span of exponentials $e^{ik\theta}$ with $0<|k|\leq N/2-1$. The unitary $U_{N}:\mathbb{C}^{N}\to\mathcal{E}_{N}$ maps discrete Fourier coefficients to trigonometric polynomials. For $L_{N}$ circulant, the lifted operator $T_{N}=U_{N}L_{N}U_{N}^{\ast}$ acts as a Fourier multiplier with discrete symbol $m_{N}$ [gray2006toeplitz].
@@
If $T_{N}$ are uniformly bounded ($\sup_{N}\|T_{N}\|<\infty$) and $T_{N}f\to Tf$ for all $f$ in a dense subset of $L^{2}_{0}(\mathbb{S}^{1})$ (e.g. trigonometric polynomials), then $T_{N}\to T$ strongly on $L^{2}_{0}(\mathbb{S}^{1})$ [kreyszig1989introductory].